Find the solution to the problem with the following initial value:

с22 dy 1 — 0, х > 0, у(0) — 1. у(1 + 2) х dx


We ask for an explicit solution. Justify each step of your solution.


Indication: В и = A + 1 u 1u
where A and B are constants.
Answer #1

Since they * dx y(1+x) So, e yudy = within So, ker (256- Cei So, ker(20= (*1kits So, Leo (25d»)=(-(ie tydens So, bye (234=So, bem =x – In}1+x+c Since y(0) =1 So, Fe = 0 – In]1+01+c So, c= Put c=in že = x= In|1+x/+C So, bem = x= 1n|1+x1+ So, eu? =